\begin{thebibliography}{10}

\bibitem{eye-tracking-DL-asd-detection-ahmed-2022}
Ibrahim~Abdulrab Ahmed, Ebrahim~Mohammed Senan, Taha~H. Rassem, Mohammed A.~H. Ali, Hamzeh Salameh~Ahmad Shatnawi, Salwa~Mutahar Alwazer, and Mohammed Alshahrani.
\newblock Eye tracking-based diagnosis and early detection of autism spectrum disorder using machine learning and deep learning techniques.
\newblock {\em Electronics}, 11(4), 2022.

\bibitem{Chen_2019_ICCV}
Shi Chen and Qi~Zhao.
\newblock Attention-based autism spectrum disorder screening with privileged modality.
\newblock In {\em Proceedings of the IEEE/CVF International Conference on Computer Vision (ICCV)}, October 2019.

\bibitem{vit-paper-2020}
Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, Jakob Uszkoreit, and Neil Houlsby.
\newblock An image is worth 16x16 words: Transformers for image recognition at scale, 2021.

\bibitem{fang-gaze-cnn-lstm-2020}
Yi~Fang, Huiyu Duan, Fangyu Shi, Xiongkuo Min, and Guangtao Zhai.
\newblock Identifying children with autism spectrum disorder based on gaze-following.
\newblock In {\em 2020 IEEE International Conference on Image Processing (ICIP)}, pages 423--427, 2020.

\bibitem{ms-celeb-dataset-paper-2016}
Yandong Guo, Lei Zhang, Yuxiao Hu, Xiaodong He, and Jianfeng Gao.
\newblock Ms-celeb-1m: A dataset and benchmark for large-scale face recognition, 2016.

\bibitem{deep-svm-meta-learning-asd-diagnosis-2023}
Xiangmin Han, Jun Wang, Shihui Ying, Jun Shi, and Dinggang Shen.
\newblock Ml-dsvm+: A meta-learning based deep svm+ for computer-aided diagnosis.
\newblock {\em Pattern Recognition}, 134:109076, 2023.

\bibitem{resnet-paper-2015}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun.
\newblock Deep residual learning for image recognition, 2015.

\bibitem{he2016deep}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun.
\newblock Deep residual learning for image recognition.
\newblock In {\em Proceedings of the IEEE conference on computer vision and pattern recognition}, pages 770--778, 2016.

\bibitem{hu-fmri-semi-supervised-lstm-asd-detection-2023}
Yao Hu, Zhi-An Huang, Rui Liu, Xiaoming Xue, Xiaoyan Sun, Linqi Song, and Kay~Chen Tan.
\newblock Source free semi-supervised transfer learning for diagnosis of mental disorders on fmri scans.
\newblock {\em IEEE Transactions on Pattern Analysis and Machine Intelligence}, 45(11):13778--13795, 2023.

\bibitem{image-asd-paper-2023}
Assil Jaby, Md~Baharul Islam, and Md~Atiqur~Rahman Ahad.
\newblock Asd-evnet: An ensemble vision network based on facial expression for autism spectrum disorder recognition.
\newblock In {\em 2023 18th International Conference on Machine Vision and Applications (MVA)}, pages 1--5, 2023.

\bibitem{convolution-kernel-asd-detection-2021}
Junzhong Ji, Xinying Xing, Yao Yao, Junwei Li, and Xiaodan Zhang.
\newblock Convolutional kernels with an element-wise weighting mechanism for identifying abnormal brain connectivity patterns.
\newblock {\em Pattern Recognition}, 109:107570, 2021.

\bibitem{Jiang_2017_ICCV}
Ming Jiang and Qi~Zhao.
\newblock Learning visual attention to identify people with autism spectrum disorder.
\newblock In {\em Proceedings of the IEEE International Conference on Computer Vision (ICCV)}, Oct 2017.

\bibitem{swin-transformer-paper-2021}
Ze~Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and Baining Guo.
\newblock Swin transformer: Hierarchical vision transformer using shifted windows, 2021.

\bibitem{cosine-scheduler-paper}
Ilya Loshchilov and Frank Hutter.
\newblock Sgdr: Stochastic gradient descent with warm restarts, 2017.

\bibitem{adamw-optimizer-paper}
Ilya Loshchilov and Frank Hutter.
\newblock Decoupled weight decay regularization, 2019.

\bibitem{pyramid-kernel-asd-diagnosis-2023}
Kai Ma, Shuo Huang, Peng Wan, and Daoqiang Zhang.
\newblock Optimal transport based pyramid graph kernel for autism spectrum disorder diagnosis.
\newblock {\em Pattern Recognition}, 143:109716, 2023.

\bibitem{eye-track-asd-detetion-mazumdar-2021}
Pramit Mazumdar, Giuliano Arru, and Federica Battisti.
\newblock Early detection of children with autism spectrum disorder based on visual exploration of images.
\newblock {\em Signal Processing: Image Communication}, 94:116184, 2021.

\bibitem{mehta2022mobilevit}
Sachin Mehta and Mohammad Rastegari.
\newblock Mobilevit: Light-weight, general-purpose, and mobile-friendly vision transformer, 2022.

\bibitem{nguyen2023insect}
Hoang-Quan Nguyen, Thanh-Dat Truong, Xuan~Bac Nguyen, Ashley Dowling, Xin Li, and Khoa Luu.
\newblock Insect-foundation: A foundation model and large-scale 1m dataset for visual insect understanding.
\newblock {\em arXiv preprint arXiv:2311.15206}, 2023.

\bibitem{nguyen2021clusformer}
Xuan-Bac Nguyen, Duc~Toan Bui, Chi~Nhan Duong, Tien~D Bui, and Khoa Luu.
\newblock Clusformer: A transformer based clustering approach to unsupervised large-scale face and visual landmark recognition.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 10847--10856, 2021.

\bibitem{nguyen2023micron}
Xuan-Bac Nguyen, Chi~Nhan Duong, Xin Li, Susan Gauch, Han-Seok Seo, and Khoa Luu.
\newblock Micron-bert: Bert-based facial micro-expression recognition.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 1482--1492, 2023.

\bibitem{nguyen2023fairness}
Xuan-Bac Nguyen, Chi~Nhan Duong, Marios Savvides, Kaushik Roy, and Khoa Luu.
\newblock Fairness in visual clustering: A novel transformer clustering approach.
\newblock {\em arXiv preprint arXiv:2304.07408}, 2023.

\bibitem{nguyen2020self}
Xuan-Bac Nguyen, Guee~Sang Lee, Soo~Hyung Kim, and Hyung~Jeong Yang.
\newblock Self-supervised learning based on spatial awareness for medical image analysis.
\newblock {\em IEEE Access}, 8:162973--162981, 2020.

\bibitem{nguyen2023brainformer}
Xuan-Bac Nguyen, Xin Li, Samee~U Khan, and Khoa Luu.
\newblock Brainformer: Modeling mri brain functions to machine vision.
\newblock {\em arXiv preprint arXiv:2312.00236}, 2023.

\bibitem{pytorch-paper}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu~Fang, Junjie Bai, and Soumith Chintala.
\newblock Pytorch: An imperative style, high-performance deep learning library, 2019.

\bibitem{mri-asd-detection-rakic-2020}
Mladen Rakić, Mariano Cabezas, Kaisar Kushibar, Arnau Oliver, and Xavier Lladó.
\newblock Improving the detection of autism spectrum disorder by combining structural and functional mri information.
\newblock {\em NeuroImage: Clinical}, 25:102181, 2020.

\bibitem{fmri-asd-detection-cnn-2020}
Zeinab Sherkatghanad, Mohammadsadegh Akhondzadeh, Soorena Salari, Mariam Zomorodi-Moghadam, Moloud Abdar, U.~Rajendra Acharya, Reza Khosrowabadi, and Vahid Salari.
\newblock Automated detection of autism spectrum disorder using a convolutional neural network.
\newblock {\em Frontiers in Neuroscience}, 13, 2020.

\bibitem{efficientnet-paper-2020}
Mingxing Tan and Quoc~V. Le.
\newblock Efficientnet: Rethinking model scaling for convolutional neural networks, 2020.

\bibitem{truong2022direcformer}
Thanh-Dat Truong, Quoc-Huy Bui, Chi~Nhan Duong, Han-Seok Seo, Son~Lam Phung, Xin Li, and Khoa Luu.
\newblock Direcformer: A directed attention in transformer approach to robust action recognition.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 20030--20040, 2022.

\bibitem{activity-recognition-asd-detection-2021}
Peter Washington, Aaron Kline, Onur~Cezmi Mutlu, Emilie Leblanc, Cathy Hou, Nate Stockham, Kelley Paskov, Brianna Chrisman, and Dennis Wall.
\newblock Activity recognition with moving cameras and few training examples: Applications for detection of autism-related headbanging.
\newblock In {\em Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems}, CHI EA '21, New York, NY, USA, 2021. Association for Computing Machinery.

\bibitem{fmri-autoencoder-fscore-asd-detection-2023}
Jin Zhang, Fan Feng, Tianyi Han, Xiaoli Gong, and Feng Duan.
\newblock Detection of autism spectrum disorder using fmri functional connectivity with feature selection and deep learning.
\newblock {\em Cognitive Computation}, 2023.

\end{thebibliography}

\documentclass[conference]{IEEEtran}
\IEEEoverridecommandlockouts
\usepackage{cite}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{algorithmic}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{xcolor}
\usepackage[pagebackref,breaklinks,colorlinks]{hyperref}
\def\BibTeX{{\rm B\kern-.05em{\sc i\kern-.025em b}\kern-.08em
    T\kern-.1667em\lower.7ex\hbox{E}\kern-.125emX}}
\begin{document}

\def\needtocite{\textcolor{red}{[CITATIONS]}}
\newcommand{\red}[1]{\textcolor{red}{#1}}

\title{
    A Novel Dataset for Video-Based Neurodivergent Classification Leveraging Extra-Stimulatory Behavior
}

\author{\IEEEauthorblockN{Manuel Serna-Aguilera}
\IEEEauthorblockA{\textit{Dept. of EECS} \\
\textit{University of Arkansas}\\
Fayetteville, United States \\
mserna@uark.edu}
\and
\IEEEauthorblockN{Xuan Bac Nguyen}
\IEEEauthorblockA{\textit{Dept. of EECS} \\
\textit{University of Arkansas}\\
Fayetteville, United States \\
xnguyen@uark.edu}
\and
\IEEEauthorblockN{Han-Seok Seo}
\IEEEauthorblockA{\textit{Dept. of Food Science} \\
\textit{University of Arkansas}\\
Fayetteville, United States \\
hanseok@uark.edu}
\and
\IEEEauthorblockN{Khoa Luu}
\IEEEauthorblockA{\textit{Dept. of EECS} \\
\textit{University of Arkansas}\\
Fayetteville, United States \\
khoaluu@uark.edu}
}

\maketitle

% Notes:
% IEEE Greentech is not double-blind review, so we can keep the author names visible in our submission

%****************************************
% Abstract
%****************************************
\begin{abstract}
Facial expressions and actions differ among different individuals at varying degrees of intensity given responses to external stimuli, particularly among those that are neurodivergent. Such behaviors affect people in terms of overall health, communication, and sensory processing. 
Deep learning can be responsibly leveraged to improve productivity in addressing this task, and help medical professionals to accurately understand such behaviors.
In this work, we introduce the Video ASD dataset—a dataset that contains video frame convolutional and attention map feature data—to foster further progress in the task of ASD classification.
Unlike many recent studies in ASD classification with MRI data, which require expensive specialized equipment, our method utilizes a powerful but relatively affordable GPU, a standard computer setup, and a video camera for inference. Results show that our model effectively generalizes and understands key differences in the distinct movements of the children.
Additionally, we test foundation models on this data to showcase how movement noise affects performance and the need for more data and more complex labels.
\end{abstract}

%\begin{IEEEkeywords}
%Deep Learning, Facial Expressions, Video, Classification
%\end{IEEEkeywords}

%****************************************
% One-page of greentech paper
%****************************************
\textbf{Introduction and Related Work. }
Deep learning has seen success in various deep learning tasks \cite{truong2022direcformer, nguyen2023brainformer, nguyen2023fairness, he2016deep, nguyen2023insect, nguyen2021clusformer, nguyen2023micron, nguyen2020self}, but recognizing atypical facial expressions is a relatively unexplored area. This can be attributed to the lack of a sufficiently-large dataset to help researchers better understand such behaviors. 
There are current works that use images, but not video \cite{image-asd-paper-2023} but do not make use of crucial temporal information for behavior analysis with Transformer models \cite{vit-paper-2020, swin-transformer-paper-2021, mehta2022mobilevit}. Further works analyze behaviors via video \cite{activity-recognition-asd-detection-2021}, gaze analysis \cite{Jiang_2017_ICCV, fang-gaze-cnn-lstm-2020, eye-track-asd-detetion-mazumdar-2021, eye-tracking-DL-asd-detection-ahmed-2022}, other behaviors \cite{Chen_2019_ICCV}, or via expensive equipment such as fMRI \cite{pyramid-kernel-asd-diagnosis-2023, hu-fmri-semi-supervised-lstm-asd-detection-2023, fmri-autoencoder-fscore-asd-detection-2023, fmri-asd-detection-cnn-2020, mri-asd-detection-rakic-2020, deep-svm-meta-learning-asd-diagnosis-2023, convolution-kernel-asd-detection-2021}, which may differ in degree between neurodivergent individuals.
By contrast, we analyze ASD-related behaviors by explicitly evoking reactions in more controlled settings.

\textbf{Data Collection. }
%Our dataset collection has been a collaboration between The University of Arkansas and The University of Texas at San Antonio. 
Each instance of our dataset contains a sequence of frame features for each video, thus obscuring the identify of the subjects, the ASD classification label (yes or no), and a text description. The frame features describe a person reacting to a sensory stimulus (e.g., taste or smell), with the video recording the subject's reaction in terms of the head pose (behavior), and facial expression. The text descriptions provide a description of the expression for every six second interval, e.g., ``The face expression changes from neutral to disgust'' if the face expression changes within that time frame.

\textbf{Methodology. }
Our method consists of two main components. First is the encoder, which is itself composed to two sub-encoder modules. We have one model (EfficientNet B0 CNN \cite{efficientnet-paper-2020}) to focus on neurotypical behaviors while the other model (ResNet-18 \cite{resnet-paper-2015} pretrained for the facial expression recognition task on MS-Celeb \cite{ms-celeb-dataset-paper-2016}) focuses on more neurodivergent behaviors. These deep learning-based models focus on spatial features. The input consists of 32 frames for each video. The second main component is the decoder, which is implemented by a spatial transformer, e.g., a vision transformer \cite{vit-paper-2020}. We then use a fully-connected network (or MLP) to output the probabilites for the classes. 
When we sample video slices, we average the probability predictions to obtain a final prediction.

\textbf{Experimental Results. }
% Implementation and hyperparams
We implement our framework using Pytorch \cite{pytorch-paper}.
We train and test our model using five $k$ folds, with four folds used as training samples, and the last fold for testing.
Our model trains with the AdamW optimizer \cite{adamw-optimizer-paper} with a cosine annealing scheduler \cite{cosine-scheduler-paper}, a learning rate 0.0001, and a cross entropy loss. We use a batch size of 4 on a Quadro RTX 8000 GPU with 48 GB of memory. As a preprocessing step, we perform face alignment via landmark detection and pose angle estimation is used to filter out bad frames. All input images have the spatial dimensions of $224 \times 224$.
For evaluation, we compute the classification accuracy at the per-video level.
The test accuracy on the test fold is 81.48\% with an F1 score of 0.7289. The result indicates the model is able to generalize well to similar but unseen samples, despite limited data and the number of frames processed per video. Despite this, such a low number frames may hinder the generalizability of the model, and movement noise of the subjects may contribute significantly to performance. Hence, we filter out frames where any head pose angle is out of the range $[-10, 10]$ to keep the data as controlled and noise-free as possible.
Future work will involve addressing how to incorporate frames with more extreme head poses, allowing for the analysis of up to hundreds of more frames with critical information, and account for other kinds of noise like movement and occlusions of the face.

%****************************************
% References
%****************************************
\bibliography{bib/egbib}{}
\bibliographystyle{plain}

\end{document}